OpenStudy (anonymous):

Grant has fraction 2/3 of a cup of ice cream in a bowl. He then removes two fraction 1/6 cups of ice cream to share with his little brother. Which of the following best explains the amount of ice cream left in Grant's bowl?

OpenStudy (anonymous):

Anyone wanna help?

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

am I reading this wrong.... he has 2/3 of a cup....and he gives bro 2 1/6 cups of that.....he is giving more then he has

OpenStudy (anonymous):

Idk....:/

OpenStudy (anonymous):

he is giving 2 1/6

OpenStudy (anonymous):

2 of them

OpenStudy (texaschic101):

but he only has 2/3...thats less then 2 1/6

OpenStudy (anonymous):

No its 1/6 but he gives him 2

OpenStudy (texaschic101):

oh....I see....hold on

OpenStudy (anonymous):

If that makes sense

OpenStudy (texaschic101):

he has 2/3.....he gives (1/6 * 2) = 1/3 2/3 - 1/3 = 1/3.....he has 1/3 of a cup left

OpenStudy (anonymous):

Oh ok thank you

OpenStudy (texaschic101):

sure thing :)

OpenStudy (anonymous):

Ok I got a problem ? fraction 2 over 6, because fraction 2 over 3 minus 2 times fraction 1 over 6 equals fraction 2 over 3 minus fraction 3 over 6 fraction 2 over 6, because fraction 2 over 3 minus 2 times fraction 1 over 6 equals fraction 2 over 3 minus fraction 2 over 6 fraction 1 over 6, because fraction 2 over 3 minus 2 times fraction 1 over 6 equals fraction 2 over 3 minus fraction 3 over 6 fraction 1 over 6, because fraction 2 over 3 minus 2 times fraction 1 over 6 equals fraction 2 over 3 minus fraction 2 over 6

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

2/6...because 2/3 - (2 * 1/6) = 2/3 - 2/6 ....its B

OpenStudy (anonymous):

Ok thanks

OpenStudy (texaschic101):

2/3 - 2/6 = 4/6 - 2/6 = 2/6 which rounds to 1/3....so basically, when rounded, he has 1/3 cup of ice cream left